### v1

# Serre duality for dg-algebras

Abstract.
We generalize Yekutieli-Zhang’s noncommutative Serre Duality Theorem to the setting of noncommutative spaces associated to dg-algebras. As an application, we establish some finiteness properties of derived global sections over such noncommutative spaces. Along the way, we generalize Yekutieli’s notion of a balanced dualizing complex to the setting of dg-algebras and establish some cases in which they exist.
Mathematics Subject Classification 2020: 14F08
1. Introduction
Let $\mathbf{k}$ be a field and $X$ a smooth and proper $\mathbf{k}$-variety of dimension $n$. Given a vector bundle $\mathcal{V}$ on $X$, the classical Serre Duality Theorem states that there are isomorphisms
(1.1) $H^{j}(X,\mathcal{V})\cong\operatorname{Ext}^{n-j}(\mathcal{V},\omega_{X})^{*}$
for all $0\leq j\leq n$, where $\omega_{X}$ denotes the canonical bundle of $X$, and the superscript * denotes the $\mathbf{k}$-dual. Serre duality has been generalized in many different directions. For instance, Grothendieck-Verdier Duality extends the statement of Serre duality to proper morphisms of varieties [Har66, Theorem 3.3], and Bondal-Kapranov’s notion of a Serre functor on a triangulated category [BK89] has led to the discovery of analogues of Serre duality for a host of categories arising in algebraic geometry, commutative algebra, and representation theory. We refer the reader to [RVdB02] for many examples of derived categories with Serre functors; we refer also to [Mur13] for a discussion of the Serre functor on the singularity category of a commutative Gorenstein local ring with isolated singularity (see also [Aus78, Buc21]).
We focus in this paper on an extension of Serre duality to graded noncommutative algebras due to Yekutieli-Zhang [YZ97]. Let $A=\bigoplus_{i\geq 0}A_{i}$ be a graded Noetherian algebra such that $A_{0}=\mathbf{k}$. Denote by $\operatorname{D}_{\operatorname{Qgr}}(A)$ the quotient of the derived category $\operatorname{D}(A)$ of graded right $A$-modules by its subcategory $\operatorname{D}^{\operatorname{Tors}}(A)$ of complexes with torsion cohomology, i.e. complexes such that each cohomology class is annihilated by some power of the homogeneous maximal ideal. Given an object $M\in\operatorname{D}(A)$, we denote its image in $\operatorname{D}_{\operatorname{Qgr}}(A)$ by $\widetilde{M}$. When $A$ is commutative and generated in degree 1, the category $\operatorname{D}_{\operatorname{Qgr}}(A)$ is equivalent to the derived category of quasicoherent sheaves on the projective scheme $X=\operatorname{Proj}(A)$. In general, no such scheme $X$ exists, but the intuition underlying the theory of noncommutative projective schemes, pioneered by Artin-Zhang [AZ94], is that $\operatorname{D}_{\operatorname{Qgr}}(A)$ nevertheless shares many features of the derived category of a projective scheme. One such feature we consider in this paper is a noncommutative analogue of the derived global sections functor [AZ94, Section 7], denoted $\mathbf{R}\Gamma_{*}\colon\operatorname{D}_{\operatorname{Qgr}}(A)\to\operatorname{D}(A)$. Given $j\in\mathbb{Z}$ and $\widetilde{M},\widetilde{N}\in\operatorname{D}_{\operatorname{Qgr}}(A)$, let $\mathbf{R}^{j}\Gamma(\widetilde{M})$ denote the internal degree 0 part of $H^{j}\mathbf{R}\Gamma_{*}(\widetilde{M})$, and let $\operatorname{Ext}^{j}(\widetilde{M},\widetilde{N})\coloneqq\operatorname{Hom}_{\operatorname{D}_{\operatorname{Qgr}}(A)}(\widetilde{M},\widetilde{N}[j])$. Yekutieli-Zhang proved the following noncommutative version of Serre duality:
Theorem 1.2 ([YZ97] Theorem 4.2(2)).
Let $A$ be as above, and assume $A$ admits a balanced dualizing complex $R$, in the sense of [Yek92, Definitions 3.3 and 4.1]. Given $j\in\mathbb{Z}$ and an object $M$ in the bounded derived category $\operatorname{D}^{\operatorname{b}}(A)$ of graded $A$-modules, there is an isomorphism
$\mathbf{R}^{j}\Gamma(\widetilde{M})\cong\operatorname{Ext}^{-j-1}(\widetilde{M},\widetilde{R})^{*},$
where the superscript * denotes the $\mathbf{k}$-dual.
For instance, it follows from [Yek92, Section 4] that $A$ admits a balanced dualizing complex when it is AS-Gorenstein; see [Yek92, Corollary 5.6 and Theorem 7.3] for additional examples. If $A$ is AS-Gorenstein, then $R=A(-a)[n]$, where $n$ is the injective dimension of $A$, and $a\in\mathbb{Z}$ is such that there is an isomorphism $\mathbf{R}\underline{\operatorname{Hom}}_{A}(\mathbf{k},A)\simeq\mathbf{k}(a)[-n]$ in $\operatorname{D}(A)$. Theorem 1.2 therefore implies $\mathbf{R}^{j}\Gamma(\widetilde{M})\cong\operatorname{Ext}^{n-1-j}(\widetilde{M},\widetilde{A(-a)})^{*}$, which extends the Serre duality isomorphism (1.1) in the case of an arithmetically Gorenstein projective variety.
Suppose now that $A$ is a differential bigraded $\mathbf{k}$-algebra (Definition 2.1)—which we abbreviate in this paper to “dg-algebra”—and assume $A$ satisfies the conditions in Setup 2.5 below. The authors observe in their previous paper [BS23] that much of Artin-Zhang’s theory of noncommutative projective schemes generalizes to such dg-algebras. In particular, letting $\operatorname{D}(A)$ denote the derived category of differential bigraded $A$-modules (see Section 2), and defining $\operatorname{D}_{\operatorname{Qgr}}(A)$ as the quotient of $\operatorname{D}(A)$ by objects whose cohomology is annihilated by some power of the homogeneous maximal ideal of $H^{0}(A)$, we have a derived global sections functor $\mathbf{R}\Gamma_{*}\colon\operatorname{D}_{\operatorname{Qgr}}(A)\to\operatorname{D}(A);$ see [BS23, Section 3.1] or Section 4.1 below. As above, we denote the image of $M\in\operatorname{D}(A)$ in $\operatorname{D}_{\operatorname{Qgr}}(A)$ by $\widetilde{M}$. Given $\widetilde{M},\widetilde{N}\in\operatorname{D}_{\operatorname{Qgr}}(A)$ and $j\in\mathbb{Z}$, we define $\mathbf{R}^{j}\Gamma(\widetilde{M})$ and $\operatorname{Ext}^{j}(\widetilde{M},\widetilde{N})$ just as before. Our main result is a generalization of Yekutieli-Zhang’s noncommutative Serre Duality Theorem (Theorem 1.2) to dg-algebras:
Theorem 1.3.
Let $A$ be a differential bigraded $\mathbf{k}$-algebra as in Setup 2.5, and assume $A$ admits a balanced dualizing dg-module $R$ (Definition 3.13). Given $j\in\mathbb{Z}$ and an object $M$ in the bounded derived category $\operatorname{D}^{\operatorname{b}}(A)$ of differential bigraded $A$-modules (see Section 2.2), there is an isomorphism
$\mathbf{R}^{j}\Gamma(\widetilde{M})\cong\operatorname{Ext}^{-j-1}(\widetilde{M},\widetilde{R})^{*},$
where the superscript * denotes the $\mathbf{k}$-dual.
Our notion of a balanced dualizing dg-module is a generalization of [Yek92, Definition 4.1]. We prove in Theorem 3.18 that any dg-algebra $A$ as in Setup 2.5 that is Gorenstein (Definition 2.7) and graded commutative admits a balanced dualizing dg-module; we refer the reader to [BS23, Examples 2.23—2.26] for several families of examples of such dg-algebras. For instance, Koszul complexes on sequences of homogeneous forms in graded Gorenstein rings admit balanced dualizing dg-modules. We also prove in Proposition 3.15 that if $A$ is a dg-algebra as in Setup 2.5, and $\dim_{\mathbf{k}}H(A)<\infty$, then $A$ admits a balanced dualizing dg-module.
The existence of balanced dualizing dg-modules for graded commutative Gorenstein dg-algebras (Theorem 3.18) is the most technically challenging result in this paper; it is an application of the theory of minimal injective resolutions of dg-modules recently developed by Minamoto [Min21a, Min21b]. See Remark 3.21 for why the graded commutativity assumption is necessary for our proof of Theorem 3.18. We conjecture that the graded commutativity assumption may be removed from Theorem 3.18: see Conjecture 3.22.
As an application of our Serre duality result (Theorem 1.3), we prove that, if $A$ is a Gorenstein dg-algebra that admits a balanced dualizing dg-module, and $M\in\operatorname{D}^{\operatorname{b}}(A)$, then the truncation $\mathbf{R}\Gamma(\widetilde{M})_{\geq i}$ of the derived global sections of $\widetilde{M}$ in internal degrees at least $i$ is an object in $\operatorname{D}^{\operatorname{b}}(A)$. That is, $\mathbf{R}\Gamma(\widetilde{M})_{\geq i}$ has bounded cohomology, and each $\mathbf{R}^{j}\Gamma(\widetilde{M})_{\geq i}$ is finitely generated over $H^{0}(A)$: see Theorem 5.1 for the precise statement. As a consequence, we extend the authors’ dg-algebra version of Orlov’s Landau-Ginzburg/Calabi-Yau correspondence [BS23, Theorem 1.3] to the case where $A$ is Gorenstein and admits a balanced dualizing dg-module: see Theorem 5.6. We emphasize that a proof of Conjecture 3.22 would allow one to remove from Theorems 5.1 and 5.6 the assumption that $A$ admits a balanced dualizing dg-module.
Overview of the paper.
We recall in Section 2 some background on differential bigraded algebras and their derived categories. In Section 3, we introduce a notion of local cohomology for dg-algebras $A$ as in Setup 2.5, as well as our notion of a balanced dualizing dg-module. We prove a Local Duality result for dg-$A$-modules (Theorem 3.11), and we show that balanced dualizing dg-modules exist in some cases (Proposition 3.15 and Theorem 3.18). We prove Serre duality for dg-algebras (Theorem 1.3) in Section 4. Finally, in Section 5, we discuss an application of Theorem 1.3 to finiteness properties of derived global sections (Theorem 5.1), as well as an extension of the dg-algebra version of Orlov’s Landau-Ginzburg/Calabi-Yau correspondence [BS23, Theorem 1.3] (see Theorem 5.6).
2. Preliminaries
We begin with some notation and conventions. Throughout, $\mathbf{k}$ denotes a field. We will consider bigraded $\mathbf{k}$-vector spaces $V=\bigoplus_{i,j\in\mathbb{Z}}V_{i}^{j}$, where the superscript typically denotes a cohomological grading, and the subscript refers to an internal grading. Given $v\in V^{j}_{i}$, we write $\deg(v)=i$ and $|v|=j$. We will denote the $m^{\operatorname{th}}$ shift of $V$ in internal (resp. cohomological) degree by $V(m)$ (resp. $V[m]$). That is, $V(m)_{i}^{j}=V_{i+m}^{j}$, and $V[m]_{i}^{j}=V_{i}^{j+m}$.
2.1. Differential bigraded algebras and modules
Definition 2.1.
A differential bigraded $\mathbf{k}$-algebra is a bigraded $\mathbf{k}$-algebra $A=\bigoplus_{(i,j)\in\mathbb{Z}^{2}}A_{i}^{j}$ equipped with a degree $(0,1)$ $\mathbf{k}$-linear map $\partial_{A}$ that squares to 0 and satisfies the Leibniz rule:
$\partial_{A}(xy)=\partial_{A}(x)y+(-1)^{|x|}x\partial_{A}(y).$
In this paper, we will call differential bigraded $\mathbf{k}$-algebras dg-algebras, for short. We say a dg-algebra $A$ is graded commutative if $xy=(-1)^{|x||y|}yx$ for all homogeneous $x,y\in A$. We will simply call $A$ commutative if it is commutative in the usual sense, i.e. $xy=yx$ for all $x,y\in A$.
Let $A$ be a dg-algebra. We denote the underlying bigraded $k$-algebra of $A$ by $A^{\natural}$. A right (resp. left) dg-$A$-module is a right (resp. left) bigraded $A^{\natural}$-module $M=\bigoplus_{i,j\in\mathbb{Z}}M^{j}_{i}$ equipped with a bidegree $(0,1)$ differential that satisfies the Leibniz rule: see [BS23, Definition 2.2] for details. All modules are assumed to be right modules unless otherwise noted. Morphisms and quasi-isomorphisms of dg-algebras and dg-modules are defined in the evident way: once again, we refer to [BS23, Definitions 2.1 and 2.2] for details. We let $A^{\operatorname{op}}$ denote the opposite dg-algebra of $A$, as defined in [BS23, Definition 2.1]. If $A$ and $B$ are dg-algebras, an $A$-$B$-bimodule is a right $A^{\operatorname{op}}\otimes_{k}B$ bimodule.
Given a dg-$A$-module $M$, we set $M_{i}\coloneqq\bigoplus_{j\in\mathbb{Z}}M_{i}^{j}$ for all $i\in\mathbb{Z}$. We observe that $A_{0}$ is a dg-algebra (concentrated in internal degree 0), and each $M_{i}$ is a dg-$A_{0}$-module (concentrated in internal degree $i$). Similarly, we set $M^{j}\coloneqq\bigoplus_{i\in\mathbb{Z}}M^{j}_{i}$ for all $j\in\mathbb{Z}$. We say a dg-algebra $A$ is connected if $A_{0}=A_{0}^{0}=\mathbf{k}$, and $A_{i}^{j}=0$ when $i<0$ or $j>0$. Notice that, if $A$ is connected, then $A_{0}=\mathbf{k}$ is a dg-$A$-module.
We recall the definitions of the tensor product and internal $\operatorname{Hom}$ for dg-$A$-modules, as described, for instance, in [BS23, Section 2]. Given a dg-algebra $A$, a right dg-$A$-module $M$, and a left dg-$A$-module $N$, the tensor product $M\otimes_{A}N$ is a dg-$\mathbf{k}$-module with differential
$m\otimes n\mapsto d_{M}(m)\otimes n+(-1)^{|m|}m\otimes d_{N}(n).$
If $M$ (resp. $N$) is an $A$-$A$-bimodule, then $M\otimes_{A}N$ is a left (resp. right) dg-$A$-module. If $M$ and $N$ are right dg-$A$-modules, we let $\operatorname{Hom}_{A}(M,N)$ denote the $\mathbf{k}$-vector space of dg-$A$-module morphisms from $M$ to $N$. We denote by $\underline{\operatorname{Hom}}_{A}(M,N)$ the internal $\operatorname{Hom}$ from $M$ to $N$, i.e. the dg-$\mathbf{k}$-vector space with underlying bigraded $\mathbf{k}$-vector space given by the $A^{\natural}$-linear internal $\operatorname{Hom}$ space $\underline{\operatorname{Hom}}_{A^{\natural}}(M,N)$ and differential
$\alpha\mapsto d_{N}\alpha-(-1)^{|\alpha|}\alpha d_{M}.$
Notice that $\operatorname{Hom}_{A}(M,N)$ is the space of bidegree $(0,0)$ cocycles in $\underline{\operatorname{Hom}}_{A}(M,N)$. If $M$ (resp. $N$) is an $A$-$A$-bimodule, then $\underline{\operatorname{Hom}}_{A}(M,N)$ is a right (resp. left) dg-$A$-module.
2.2. Derived categories of dg-algebras
We let $\operatorname{Mod}(A)$ denote the category of dg-$A$-modules, $\operatorname{D}(A)$ the derived category of dg-$A$-modules, and $\operatorname{D}^{\operatorname{b}}(A)\subseteq\operatorname{D}(A)$ the subcategory given by objects isomorphic to dg-$A$-modules that are finitely generated over $A$. We refer the reader to [BS23, Definition 2.13] for the definitions of $K$-projective, $K$-flat, and $K$-injective resolutions of dg-modules. Given a pair $M$ and $N$ of dg-modules over a dg-algebra $A$, $K$-flat resolutions $F^{M}$ and $F^{N}$ of $M$ and $N$, a $K$-projective resolution $P$ of $M$, and a $K$-injective resolution $I$ of $N$, the derived tensor product of $M$ and $N$ and derived Hom from $M$ to $N$ may be modeled as follows:
$M\otimes_{A}^{\mathbf{L}}N\cong F^{M}\otimes_{A}N\cong M\otimes_{A}F^{N},\quad\mathbf{R}\underline{\operatorname{Hom}}_{A}(M,N)\cong\underline{\operatorname{Hom}}_{A}(P,N)\cong\underline{\operatorname{Hom}}_{A}(M,I);$
where the isomorphisms are in the derived category $\operatorname{D}(A)$. In particular, we write:
$\underline{\operatorname{Ext}}^{i}_{A}(M,N)\coloneqq H^{i}\mathbf{R}\underline{\operatorname{Hom}}_{A}(M,N)\quad\text{and}\quad\operatorname{Ext}^{i}_{A}(M,N)\coloneqq\underline{\operatorname{Ext}}^{i}_{A}(M,N)_{0}.$
Remark 2.2.
Let $K(A)$ denote the homotopy category of dg-$A$-modules. For any dg-$A$-module $M$ and $K$-injective dg-$A$-module $I$, the natural map $\operatorname{Hom}_{K(A)}(M,I)\to\operatorname{Hom}_{\operatorname{D}(A)}(M,I)$ is an isomorphism [Yek20, Theorem 10.1.13]. Similarly, if $N$ is a dg-$A$-module, and $P$ is a $K$-projective dg-$A$-module, then the natural map $\operatorname{Hom}_{K(A)}(P,N)\to\operatorname{Hom}_{\operatorname{D}(A)}(P,N)$ is an isomorphism [Yek20, Theorem 10.2.9].
Proposition 2.3.
If $M$ and $N$ are dg-$A$-$A$-bimodules, then $\mathbf{R}\underline{\operatorname{Hom}}_{A}(M,N)$ and $\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(M,N)$ may be equipped with the structure of dg-$A$-$A$-bimodules.
To prove Proposition 2.3, we will need the following technical result.
Lemma 2.4.
Let $A$ be a dg-algebra.
  1. (1)
    Let $f:A\to B$ be a dg-algebra morphism such that $B$ is $K$-flat as a left dg-$A$-module. If $I$ is a $K$-injective right dg-$B$-module, then it is also a $K$-injective right dg-$A$-module.
  2. (2)
    Let $N$ be a dg-$A$-$A$-bimodule, and let $I$ be a $K$-injective resolution of $N$ as an $A^{\operatorname{op}}\otimes_{\mathbf{k}}A$-module. The resolution $I$ is $K$-injective as both a right dg-$A$-module and a right dg-$A^{\operatorname{op}}$-module. Thus, $I$ is a $K$-injective resolution of $N$ as both a right dg-$A$-module and a right dg-$A^{\operatorname{op}}$-module.
Proof.
To prove (1), let $M$ be an exact right dg-$A$-module. Since $B$ is a $K$-flat left dg-$A$-module, the right dg-$B$-module $M\otimes_{A}B$ is exact. It follows that $\operatorname{\underline{Hom}}_{B}(M\otimes_{A}B,I)$ is exact. By adjunction, we have $\operatorname{\underline{Hom}}_{B}(M\otimes_{A}B,I)\cong\operatorname{\underline{Hom}}_{A}(M,\operatorname{\underline{Hom}}_{B}(B,I))\cong\operatorname{\underline{Hom}}_{A}(M,I).$ This proves (1). As for (2): since $A^{\operatorname{op}}\otimes_{\mathbf{k}}A$ is $K$-flat as both a left dg-$A$-module and a left dg-$A^{\operatorname{op}}$-module, Lemma 2.4 implies that $I$ is $K$-injective as both a right dg-$A$-module and a right dg-$A^{\operatorname{op}}$-module. ∎
Proof of Proposition 2.3.
Choose a $K$-injective resolution $I$ of $N$ as a dg-$A^{\operatorname{op}}\otimes_{\mathbf{k}}A$-module. By Lemma 2.4(2), we see that $\mathbf{R}\underline{\operatorname{Hom}}_{A}(M,N)$ can be modeled as $\underline{\operatorname{Hom}}_{A}(M,I)$, and $\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(M,N)$ can be modeled as $\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(M,I)$. It follows that both $\mathbf{R}\underline{\operatorname{Hom}}_{A}(M,N)$ and $\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(M,N)$ may be equipped with dg-$A$-$A$-bimodule structures. ∎
2.3. Gorenstein dg-algebras
For the rest of the paper, we will work under the following setup, which is identical to [BS23, Setup 2.8]:
Setup 2.5.
Let $A$ be a connected dg-algebra such that the underlying ring $A^{\natural}$ is Noetherian, $H^{0}(A)$ is Noetherian, and the total cohomology algebra $H(A)$ is finitely generated as an $H^{0}(A)$-module.
Remarks 2.6.
Let $A$ be a dg-algebra as in Setup 2.5.
  1. (1)
    Let $\operatorname{D}^{\operatorname{f}}(A)$ denote the subcategory of $\operatorname{D}(A)$ given by dg-$A$-modules $M$ such that the total cohomology $H(M)$ is finitely generated over the total cohomology $H(A)$. By [BS23, Proposition 2.17], the canonical map $\operatorname{D}^{\operatorname{b}}(A)\to\operatorname{D}^{\operatorname{f}}(A)$ is an equivalence.
  2. (2)
    If $M\in\operatorname{D}^{\operatorname{b}}(A)$, then it follows from [BS23, Proposition 2.16] that $\dim_{\mathbf{k}}H(M)_{\ell}<\infty$ for all $\ell\in\mathbb{Z}$.
Definition 2.7.
Let $A$ be as in Setup 2.5. We say $A$ is Gorenstein if:
  1. (1)
    The functor $\mathbf{R}\underline{\operatorname{Hom}}_{A}(-,A)$ maps $\operatorname{D}^{\operatorname{b}}(A)$ to $\operatorname{D}^{\operatorname{b}}(A^{\operatorname{op}})$, and $\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(-,A)$ maps $\operatorname{D}^{\operatorname{b}}(A^{\operatorname{op}})$ to $\operatorname{D}^{\operatorname{b}}(A)$.
  2. (2)
    Given $M\in\operatorname{D}^{\operatorname{b}}(A)$ and $N\in\operatorname{D}^{\operatorname{b}}(A^{\operatorname{op}})$, the canonical maps
    $M\to\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(\mathbf{R}\underline{\operatorname{Hom}}_{A}(M,A),A)\quad\text{and}\quad N\to\mathbf{R}\underline{\operatorname{Hom}}_{A}(\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(N,A),A)$
    are isomorphisms in $\operatorname{D}(A)$ and $\operatorname{D}(A^{\operatorname{op}})$ respectively.
  3. (3)
    There is an isomorphism $\mathbf{R}\underline{\operatorname{Hom}}_{A}(\mathbf{k},A)\cong\mathbf{k}(a)[-n]$ in $\operatorname{D}(A)$ for some $a,n\in\mathbb{Z}$.
When $A$ is Gorenstein, the integer $a$ in (3) is called the Gorenstein parameter of $A$.
There are other definitions of Gorenstein dg-algebras in the literature: see [BS23, Remark 2.16] for a detailed discussion. We refer to [BS23, Examples 2.23—2.27] for several examples of Gorenstein dg-algebras.
2.4. Minamoto’s dg version of minimal injective resolutions
Let $A$ be as in Setup 2.5. We now discuss (a bigraded variant of) a notion of minimal injective resolutions for dg-modules recently developed by Minamoto [Min21a, Min21b]. We first observe that, since $A$ is connected, the inclusion $A^{0}\to A$ is a morphism of dg-algebras; in particular, $A$ is a dg-$A^{0}$-module. Moreover, the complex $\operatorname{\underline{Hom}}_{A^{0}}(A,M)$ is naturally a dg-$A$-module for any dg-$A^{0}$-module $M$.
Notation 2.8.
([Min21b] Definition 3.9) Let $\mathrm{Inj}(A^{0})$ denote the category of graded injective $A^{0}$-modules (not dg-modules), and let $\mathcal{I}$ be the full subcategory of $\operatorname{D}(A)$ given by objects isomorphic to $\operatorname{\underline{Hom}}_{A^{0}}(A,K)$ for some $K\in\mathrm{Inj}(A^{0})$.
Remark 2.9.
Any object in $\mathcal{I}$ is both $K$-injective and injective as an $A^{\natural}$-module [Min21b, Corollary 3.13].
The following is a bigraded analogue of [Min21b, Theorem 3.10]; its proof is essentially the same.
Theorem 2.10 (cf. [Min21b] Theorem 3.10).
Let $A$ be as in Setup 2.5.
  1. (1)
    If $M\in\operatorname{D}(A)$, and $I\in\mathcal{I}$, then $\operatorname{Hom}_{\operatorname{D}(A)}(M,I)\cong\operatorname{Hom}_{H^{0}(A)}(H^{0}(M),H^{0}(I))$.
  2. (2)
    The 00-th cohomology functor $H^{0}:\operatorname{D}(A)\rightarrow\operatorname{Mod}(H^{0}(A))$ induces an equivalence
    $H^{0}:\mathcal{I}\xrightarrow{\simeq}\mathrm{Inj}(H^{0}(A)).$
Proof.
Let $M\in\operatorname{D}(A)$ and $I\in\mathcal{I}$. Choose $K\in\mathrm{Inj}(A^{0})$ such that $I\cong\operatorname{\underline{Hom}}_{A^{0}}(A,K)$. We note that:
(2.11) $\operatorname{\underline{Hom}}_{A^{0}}(H^{0}(N),K)\cong H^{0}(\operatorname{\underline{Hom}}_{A^{0}}(N,K))\quad\text{for any dg-$A$-module $N$ and any injective $A^{0}$-module $K$}.$
We have:
$\operatorname{Hom}_{\operatorname{D}(A)}(M,I)\cong\operatorname{Hom}_{K(A)}(M,I)=H^{0}(\operatorname{\underline{Hom}}_{A}(M,I))_{0}$ $\cong H^{0}(\operatorname{\underline{Hom}}_{A^{0}}(M,K))_{0}$
$\cong\operatorname{Hom}_{A^{0}}(H^{0}(M),K)$
$\cong\operatorname{Hom}_{A^{0}}(H^{0}(M),\operatorname{\underline{Hom}}_{A^{0}}(H^{0}(A),K)),$
where the first isomorphism follows from Remark 2.2, the second from adjunction, the third from (2.11), and the last from adjunction again, along with the observation that $H^{0}(M)\otimes_{A^{0}}H^{0}(A)\cong H^{0}(M)$. Applying (2.11) once more proves (1). For (2), we first observe that (1) implies that the functor $H^{0}\colon\mathcal{I}\to\operatorname{Inj}(H^{0}(A))$ is fully faithful. A bigraded analogue of [Min21b, Lemma 3.14] implies that every injective $H^{0}(A)$-module is of the form $\underline{\operatorname{Hom}}_{A^{0}}(H^{0}(A),K)$ for some injective $A^{0}$-module $K$; applying (2.11) therefore finishes the proof. ∎
We now give a bigraded version of Minamoto’s notion of minimal injective resolutions of dg-modules, which Minamoto calls “minimal ifij resolutions” in [Min21b]. Before stating it, we fix some notation: given a dg-$A$-module $M$, we write
(2.12) $\sup(M)=\sup\{i\text{ : }H^{i}(M)\neq 0\}\quad\text{and}\quad\inf(M)=\inf\{i\text{ : }H^{i}(M)\neq 0\}.$
Definition 2.13 ([Min21b] Definition 3.19).
Let $A$ be as in Setup 2.5, and let $M$ be a dg-$A$-module such that $M$ is not exact and $H^{i}(M)=0$ for $i\ll 0$. A minimal injective resolution of $M$ is a sequence of exact triangles $\left[M_{i}\xrightarrow{f_{i}}I_{i}\xrightarrow{g_{i}}M_{i-1}\to\right]$ in $\operatorname{D}(A)$ for $i\leq 0$ such that $M_{0}=M$, and for all $i\leq 0$:
  1. (1)
    We have $I_{i}[\inf(I_{i})]\in\mathcal{I}$, and $\inf(I_{i})=\inf(M_{i})$.
  2. (2)
    The map $f_{i}\colon H^{\inf(M_{i})}(M_{i})\to H^{\inf(M_{i})}(I_{i})$ is a graded $H^{0}(A)$-linear injective envelope.
Since every graded module over an ordinary graded algebra admits an injective hull, it follows from 2.10 that every object $M$ as in Definition 2.13 admits a minimal injective resolution.
Remark 2.14.
Let $A$ and $M$ be as in Definition 2.13, and let $\left[M_{i}\xrightarrow{f_{i}}I_{i}\xrightarrow{g_{i}}M_{i-1}\to\right]$ be a minimal injective resolution of $M$. It is observed in [Min21b, Definition 3.19] that $\inf(I_{i})\geq\inf(M)$ for all $i\leq 0$.
3. Balanced dualizing dg-modules
Yekutieli introduced in [Yek92] the notion of a balanced dualizing complex over a noncommutative graded algebra. The main goal of this section is to introduce a generalization of balanced dualizing complexes in the setting of dg-algebras and to establish some cases in which they exist. We also introduce a notion of local cohomology for differential bigraded algebras; see also work of Shaul [Sha18, Sha19] for a similar dg-version of local cohomology. In addition, we prove a dg-version of the Local Duality Theorem, which is a generalization of Yekutieli’s noncommutative Local Duality Theorem [Yek92, Theorem 4.18].
3.1. Local cohomology of dg-modules
Let $A$ be as in Setup 2.5 and $\mathfrak{m}$ its homogeneous maximal ideal. Given $M\in\operatorname{D}(A)$, we define $\mathbf{R}\Gamma_{\mathfrak{m}}(M)\coloneqq\underset{d\to\infty}{\operatorname{colim}}\text{ }\mathbf{R}\underline{\operatorname{Hom}}_{A}(A/A_{\geq d},M)$, the derived $\mathfrak{m}$-torsion module of $M$. The $i^{\operatorname{th}}$ cohomology module $H^{i}_{\mathfrak{m}}(M)\coloneqq\underset{d\to\infty}{\operatorname{colim}}\text{ }\underline{\operatorname{Ext}}^{i}_{A}(A/A_{\geq d},M)$ of $\mathbf{R}\Gamma_{\mathfrak{m}}(M)$ is called the $i^{\operatorname{th}}$ local cohomology of $M$ (with respect to $\mathfrak{m}$).
Let $\operatorname{D}^{\operatorname{Tors}}(A)$ be the full subcategory of $\operatorname{D}(A)$ given by objects $T$ such that every class $t\in H(T)$ satisfies $t\cdot H^{0}(A)_{\geq n}=0$ for $n\gg 0$. We let $\operatorname{D}^{\operatorname{tors}}(A)$ be the full subcategory of $\operatorname{D}^{\operatorname{b}}(A)$ given by objects that are also in $\operatorname{D}^{\operatorname{Tors}}(A)$. The functor $\mathbf{R}\Gamma_{\mathfrak{m}}$ clearly takes values in $\operatorname{D}^{\operatorname{Tors}}(A)$.
Proposition 3.1.
The functor $\mathbf{R}\Gamma_{\mathfrak{m}}$ is the right adjoint of the inclusion $\operatorname{D}^{\operatorname{Tors}}(A)\hookrightarrow\operatorname{D}(A)$.
Proof.
Let $M\in\operatorname{D}(A)$. The surjections $A\to A/A_{\geq d}$ induce maps $\mathbf{R}\underline{\operatorname{Hom}}_{A}(A/A_{\geq d},M)\to M$ for all $d\geq 0$; passing to the colimit yields a natural morphism
(3.2) $\mathbf{R}\Gamma_{\mathfrak{m}}(M)\to M.$
For any $T\in\operatorname{D}^{\operatorname{Tors}}(A)$, the morphism (3.2) induces a natural map
(3.3) $\operatorname{Hom}_{\operatorname{D}(A)}(T,\mathbf{R}\Gamma_{\mathfrak{m}}(M))\to\operatorname{Hom}_{\operatorname{D}(A)}(T,M).$
It suffices to show that  (3.3) is an isomorphism. Since $T$ is a filtered colimit of objects in $\operatorname{D}^{\operatorname{tors}}(A)$, we may assume $T\in\operatorname{D}^{\operatorname{tors}}(A)$. Moreover, since $\operatorname{D}^{\operatorname{tors}}(A)$ is generated by $\mathbf{k}(i)$ for all $i\in\mathbb{Z}$ [BS23, Proposition 3.1], we may further reduce to the case where $T=\mathbf{k}$. We now define an inverse $\operatorname{Hom}_{\operatorname{D}(A)}(\mathbf{k},M)\to\operatorname{Hom}_{\operatorname{D}(A)}(\mathbf{k},\mathbf{R}\Gamma_{\mathfrak{m}}(M))$ of the map (3.3). Choose a $K$-injective resolution $I$ of $M$. Let $\varphi\colon\underline{\operatorname{Hom}}_{A}(\mathbf{k},I)\to\underset{d\to\infty}{\operatorname{colim}}\text{ }\underline{\operatorname{Hom}}_{A}(A/A_{\geq d},I)$ denote the canonical map. Since the image of $\varphi$ is annihilated by $\mathfrak{m}$, $\varphi$ determines a chain map $\underline{\operatorname{Hom}}_{A}(\mathbf{k},I)\to\underline{\operatorname{Hom}}_{A}(\mathbf{k},\underset{d\to\infty}{\operatorname{colim}}\text{ }\underline{\operatorname{Hom}}_{A}(A/A_{\geq d},I)),$ which yields a map $\operatorname{Hom}_{K(A)}(\mathbf{k},I)\to\operatorname{Hom}_{K(A)}(\mathbf{k},\underset{d\to\infty}{\operatorname{colim}}\text{ }\underline{\operatorname{Hom}}_{A}(A/A_{\geq d},I)),$ and hence, via Remark 2.2, a map
(3.4) $\operatorname{Hom}_{\operatorname{D}(A)}(\mathbf{k},M)\to\operatorname{Hom}_{\operatorname{D}(A)}(\mathbf{k},\mathbf{R}\Gamma_{\mathfrak{m}}(M)).$
The map (3.4) gives an inverse to (3.3) when $T=\mathbf{k}$. ∎
Remark 3.5.
It follows from the proof of Proposition 3.1 that the map (3.2) is the counit of the adjunction between the inclusion $\operatorname{D}^{\operatorname{Tors}}(A)\hookrightarrow\operatorname{D}(A)$ and $\mathbf{R}\Gamma_{\mathfrak{m}}$.
If $M\in\operatorname{D}(A^{\operatorname{op}})$, denote its derived $\mathfrak{m}$-torsion module and local cohomology by $\mathbf{R}\Gamma_{\mathfrak{m}^{\operatorname{op}}}(M)$ and $H^{*}_{\mathfrak{m}^{\operatorname{op}}}(M)$.
Remark 3.6.
Let $M$ be a dg-$A$-$A$-bimodule. Applying Proposition 2.3, we may choose a $K$-injective resolution of $M$ as an $A^{\operatorname{op}}\otimes_{\mathbf{k}}A$-module that is also $K$-injective as a dg-$A$-module and a dg-$A^{\operatorname{op}}$-module. For all $d\in\mathbb{Z}$, the objects $\mathbf{R}\underline{\operatorname{Hom}}_{A}(A/A_{\geq d},M)$ and $\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(A^{\operatorname{op}}/A^{\operatorname{op}}_{\geq d},M)$ are dg-$A$-$A$-bimodules; it follows that both $\mathbf{R}\Gamma_{\mathfrak{m}}(M)$ and $\mathbf{R}\Gamma_{\mathfrak{m}^{\operatorname{op}}}(M)$ are dg-$A$-$A$-bimodules as well. Moreover, the counit map (3.2) is a morphism of dg-$A$-$A$-bimodules.
The following result plays no role in what follows, but we include it as it may be of independent interest:
Proposition 3.7.
Let $A$ be as in Setup 2.5, $\mathfrak{m}$ its homogeneous maximal ideal, and $\mathfrak{n}$ the homogeneous maximal ideal of $A^{0}$. Given a dg-$A$-module $M$, there is an isomorphism $\mathbf{R}\Gamma_{\mathfrak{m}}(M)\cong\mathbf{R}\Gamma_{\mathfrak{n}}(M)$ in $\operatorname{D}(A^{0})$.
Proof.
The proof of [BS23, Proposition 3.6] implies that the extension of scalars functor $F\colon\operatorname{D}(A^{0})\to\operatorname{D}(A)$ sends $\operatorname{D}^{\operatorname{Tors}}(A^{0})$ to $\operatorname{D}^{\operatorname{Tors}}(A)$. We therefore have a commutative square
${\operatorname{D}^{\operatorname{Tors}}(A^{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F}$${\operatorname{D}(A^{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F}$${\operatorname{D}^{\operatorname{Tors}}(A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{D}(A),}$
where the horizontal maps are inclusions. Of course, $F$ is the left adjoint of the restriction of scalars functor $G$, and it follows from Proposition 3.1 that the horizontal maps are left adjoints of the derived $\mathfrak{n}$-torsion and $\mathfrak{m}$-torsion functors; passing to right adjoints, we thus have a commutative square
[…]
3.2. Local Duality
We now prove a dg-version of Local Duality, which is a generalization of Yekutieli’s noncommutative Local Duality Theorem [Yek92, Theorem 4.18]. We begin with a bigraded version of Frankild-Iyengar-Jørgensen’s notion of a dualizing dg-module. Before stating it, we recall that, given a dg-algebra $A$ and a dg-$A$-$A$-bimodule $R$, a dg-$A$-module $M$ is called $R$-reflexive if the natural map $M\to\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(\mathbf{R}\underline{\operatorname{Hom}}_{A}(M,R),R)$ is a quasi-isomorphism.
Definition 3.8 (cf. [FIJ03] Section 1.8).
Let $A$ be as in Setup 2.5. A dg-$A$-$A$-bimodule $R$ is called dualizing if, for any $M\in\operatorname{D}^{\operatorname{b}}(A^{\operatorname{op}})$ and $N\in\operatorname{D}^{\operatorname{b}}(A)$, the following conditions are satisfied:
  1. (1)
    We have $\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(M,R)\in\operatorname{D}^{\operatorname{b}}(A)$, and $\mathbf{R}\underline{\operatorname{Hom}}_{A}(N,R)\in\operatorname{D}^{\operatorname{b}}(A^{\operatorname{op}})$.
  2. (2)
    The dg-$A$-modules $N$ and $N\otimes_{A}R$ and dg-$A^{\operatorname{op}}$-modules $M$ and $M\otimes_{A^{\operatorname{op}}}R$ are $R$-reflexive.
Remark 3.9.
The definition of a dualizing dg-module in [FIJ03, Section 1.8] requires the existence of a biprojective and biinjective resolution of $R$: see [FIJ03, Section 1.2] for the definitions of these objects. But such resolutions always exist in our setting: the existence of biprojective resolutions is standard since our ground ring $\mathbf{k}$ is a field (see e.g. [FIJ03, Proposition 1.3]), and the existence of biinjective resolutions is the content of Proposition 2.4(2). Additionally, the requirement in the definition of a dualizing dg-module in [FIJ03, Section 1.8] that $A$ is $R$-reflexive as both a dg-$A$-module and a dg-$A^{\operatorname{op}}$-module is superfluous in our setting: it follows from (2) in Definition 3.8.
Example 3.10.
If $A$ is Gorenstein (Definition 2.7), then $A(i)[j]$ is dualizing for all $i,j\in\mathbb{Z}$.
Let $A$ be as in Setup 2.5. If $A$ admits a dualizing dg-module $R$ (Definition 3.8), then we have the following duality functors, which are inverse equivalences:
$D$ $\colon\operatorname{D}^{\operatorname{b}}(A)\to\operatorname{D}^{\operatorname{b}}(A^{\operatorname{op}})^{\operatorname{op}},\quad D(M)=\mathbf{R}\underline{\operatorname{Hom}}_{A}(M,R);$
$D^{\operatorname{op}}$ $\colon\operatorname{D}^{\operatorname{b}}(A^{\operatorname{op}})^{\operatorname{op}}\to\operatorname{D}^{\operatorname{b}}(A),\quad D^{\operatorname{op}}(M)=\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(M,R).$
Theorem 3.11 (Local Duality, cf. [Yek92] Theorem 4.18).
Let $A$ be as in Setup 2.5, and assume $A$ admits a dualizing dg-module $R$. Let $M\in\operatorname{D}^{\operatorname{b}}(A)$. There is a canonical isomorphism in $\operatorname{D}(A)$
$\mathbf{R}\Gamma_{\mathfrak{m}}(M)\cong\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(D(M),\mathbf{R}\Gamma_{\mathfrak{m}}(R)).$
Proof.
Fix $n\geq 1$. By Remark 3.6, the canonical map $\mathbf{R}\Gamma_{\mathfrak{m}}(R)\to R$ given by (3.2) is a morphism of $A$-$A$-bimodules, and it induces a map $\alpha_{n}\colon\mathbf{R}\underline{\operatorname{Hom}}_{A}(A/A_{\geq n},\mathbf{R}\Gamma_{\mathfrak{m}}(R))\to D(A/A_{\geq n})$ of $A$-$A$-bimodules, which is a quasi-isomorphism due to the adjunction in Proposition 3.1. We therefore have a morphism $D(A/A_{\geq n})\to\mathbf{R}\Gamma_{\mathfrak{m}}(R)$ in $\operatorname{D}(A^{\operatorname{op}}\otimes_{\mathbf{k}}A)$ given by the diagram
(3.12) $D(A/A_{\geq n})\xleftarrow[\simeq]{\alpha_{n}}\mathbf{R}\underline{\operatorname{Hom}}_{A}(A/A_{\geq n},\mathbf{R}\Gamma_{\mathfrak{m}}(R))\to\mathbf{R}\Gamma_{\mathfrak{m}}(R),$
where the second map is induced by the surjection $A\twoheadrightarrow A/A_{\geq n}$. Let $f_{n}$ denote the composition
$\mathbf{R}\underline{\operatorname{Hom}}_{A}(A/A_{\geq n},M)\xrightarrow[\simeq]{\gamma_{n}}\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(D(M),D(A/A_{\geq n}))\xrightarrow{\beta_{n}}\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(D(M),\mathbf{R}\Gamma_{\mathfrak{m}}(R)),$
where $\gamma_{n}$ is the $A$-linear quasi-isomorphism induced by the equivalence $D$, and $\beta_{n}$ is induced by (3.12). Since filtered colimits are exact in $\operatorname{Mod}(A)$, the colimit of the maps $\gamma_{n}$ is an isomorphism in $\operatorname{D}(A)$. The colimit of the maps $\beta_{n}$ is also an isomorphism in $\operatorname{D}(A)$, since $D(M)\in\operatorname{D}^{\operatorname{b}}(A^{\operatorname{op}})$. We conclude that the colimit of the maps $f_{n}$ gives the desired isomorphism in $\operatorname{D}(A)$. ∎
3.3. Balanced dualizing dg-modules
The following definition is a direct analogue of Yekutieli’s notion of a balanced dualizing complex:
Definition 3.13 ([Yek92] Definition 4.1).
Let $A$ be a dg-algebra and $R$ a dg-$A$-$A$-bimodule. We say $R$ is a balanced dualizing dg-module if it is a dualizing dg-module in the sense of Definition 3.8, and there are isomorphisms $\mathbf{R}\Gamma_{\mathfrak{m}}(R)\cong\underline{\operatorname{Hom}}_{\mathbf{k}}(A,\mathbf{k})\cong\mathbf{R}\Gamma_{\mathfrak{m}^{\operatorname{op}}}(R)$ in $\operatorname{D}(A^{\operatorname{op}}\otimes_{\mathbf{k}}A)$.
We record the following consequence of Local Duality (Theorem 3.11):
Corollary 3.14.
If $A$ is as in Setup 2.5, $R$ is a balanced dualizing dg-module for $A$, and $M\in\operatorname{D}^{\operatorname{b}}(A)$, then there is an isomorphism $\mathbf{R}\Gamma_{\mathfrak{m}}(M)\cong\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(D(M),\underline{\operatorname{Hom}}_{\mathbf{k}}(A,\mathbf{k}))$ in $\operatorname{D}(A)$.
When $A$ is as in Setup 2.5, and $A=A^{0}$, our definition agrees with Yekutieli’s definition of a balanced dualizing complex [Yek92, Definition 4.1]. In this case, Yekutieli proves that balanced dualizing complexes exist for AS-regular algebras [Yek92, Corollary 4.14] (in fact, the proof of this statement only requires that the algebra is Gorenstein, in the sense of Definition 2.7), finitely generated $\mathbf{k}$-algebras that are finitely generated over their centers [Yek92, Corollary 5.6], and skew homogeneous coordinate rings [Yek92, Theorem 7.3]. We now verify the existence of a balanced dualizing dg-module in some examples.
Proposition 3.15.
Let $A$ be as in Setup 2.5, and assume $\dim_{\mathbf{k}}H(A)<\infty$. The dg-$A$-$A$-bimodule $\underline{\operatorname{Hom}}_{\mathbf{k}}(A,\mathbf{k})$ is a balanced dualizing dg-module for $A$.
Proof.
Set $R\coloneqq\underline{\operatorname{Hom}}_{\mathbf{k}}(A,\mathbf{k})$, and let $M\in\operatorname{D}^{\operatorname{b}}(A^{\operatorname{op}})$ and $N\in\operatorname{D}^{\operatorname{b}}(A)$. It follows by adjunction that $\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(M,R)$ has finitely generated cohomology over $H^{0}(A)$, and so it is an object in $\operatorname{D}^{\operatorname{b}}(A)$; similarly, $\mathbf{R}\underline{\operatorname{Hom}}_{A}(N,R)\in\operatorname{D}^{\operatorname{b}}(A^{\operatorname{op}})$. The natural map $N\to\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(\mathbf{R}\underline{\operatorname{Hom}}_{A}(N,R),R)$ in $\operatorname{D}^{\operatorname{b}}(A)$ is the composition of the isomorphism $N\xrightarrow{\cong}\mathbf{R}\underline{\operatorname{Hom}}_{\mathbf{k}}(\mathbf{R}\underline{\operatorname{Hom}}_{\mathbf{k}}(N,\mathbf{k}),\mathbf{k})$ and the adjunction isomorphism given by $\mathbf{R}\underline{\operatorname{Hom}}_{\mathbf{k}}(\mathbf{R}\underline{\operatorname{Hom}}_{\mathbf{k}}(N,\mathbf{k}),\mathbf{k})\xrightarrow{\cong}\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(\mathbf{R}\underline{\operatorname{Hom}}_{A}(N,R),R)$; thus $N$ is $R$-reflexive. One similarly verifies that $M$ is $R$-reflexive. Since $R\in\operatorname{D}^{\operatorname{b}}(A^{\operatorname{op}}\otimes_{k}A)$, we conclude that $N\otimes_{A}R$ and $M\otimes_{A^{\operatorname{op}}}R$ are $R$-reflexive, and so $\underline{\operatorname{Hom}}_{\mathbf{k}}(A,\mathbf{k})$ is a dualizing dg-module for $A$. Finally, the counit maps $\mathbf{R}\Gamma_{\mathfrak{m}}(R)\to R$ and $\mathbf{R}\Gamma_{\mathfrak{m}^{\operatorname{op}}}(R)\to R$ are both morphisms of dg-$A$-$A$-bimodules, and they are isomorphisms in $\operatorname{D}(A^{\operatorname{op}}\otimes_{\mathbf{k}}A)$ since $\dim_{\mathbf{k}}H(R)<\infty$. ∎
Example 3.16.
If $A$ is a (possibly noncommutative) trivial extension dg-algebra, as defined in [Jin20, Section 6], then $A$ satisfies the conditions in Setup 2.5, and $\dim_{\mathbf{k}}H(A)<\infty$. Proposition 3.15 therefore implies that $\underline{\operatorname{Hom}}_{\mathbf{k}}(A,\mathbf{k})$ is a balanced dualizing dg-module for $A$.
Remark 3.17.
Suppose $A$ is as in Proposition 3.15, and assume $A$ is also Gorenstein. Choose $a,n\in\mathbb{Z}$ such that there is an isomorphism $\mathbf{R}\underline{\operatorname{Hom}}_{A}(\mathbf{k},A)\cong\mathbf{k}(a)[-n]$ in $\operatorname{D}(A)$. We have an isomorphism of dg-$A$-$A$-bimodules:
$\underline{\operatorname{Hom}}_{\mathbf{k}}(A,\mathbf{k})\cong\underline{\operatorname{Hom}}_{\mathbf{k}}(A,\mathbf{R}\underline{\operatorname{Hom}}_{A}(\mathbf{k},A))(-a)[n]\cong A(-a)[n].$
Theorem 3.18 (cf. [Yek92] Proposition 4.4).
Suppose $A$ is graded commutative and Gorenstein (see Definition 2.7). Choose $a,n\in\mathbb{Z}$ such that $\mathbf{R}\underline{\operatorname{Hom}}_{A}(\mathbf{k},A)\cong\mathbf{k}(a)[-n]$ in $\operatorname{D}(A)$. The dg-$A$-$A$-bimodule $R\coloneqq A(-a)[n]$ is a balanced dualizing dg-module.
Proof.
As observed in Example 3.10, $R$ is a dualizing dg-module for $A$. We refer the reader to [Min21b, Definition 3.1] for the notion of injective dimension of a dg-module $N$ (which is due to Yekutieli [Yek13]). Given a dg-$A$-module $N$, we denote its injective dimension by $\operatorname{id}_{A}(N)$. By [Min21a, Theorem 3.9] [^1] Throughout this proof, we reference results in [Min21a] and [Min21b], although Minamoto does not work with dg-algebras with an internal grading. However, the bigraded versions of the referenced results hold verbatim, by the same proofs. , the injective dimension of $A$ over itself is the Krull dimension of $H^{0}(A)$. We thus have $\operatorname{id}_{A}(R)=\operatorname{id}_{A}(A)=\dim H^{0}(A)<\infty$. Given a homogeneous prime ideal $\mathfrak{p}$ in $H^{0}(A)$ and a dg-$A$-module $N$, there is a notion of Bass numbers $\mu^{n}(\mathfrak{p},N)$ due to Minamoto; we refer to [Min21a, Section 2.3.2] for the definition. Combining [Min21a, Lemma 2.26] and [Min21a, Proposition 2.29], we have that $\mu^{n}(\mathfrak{p},R)=0$ for $n\gg 0$. It therefore follows from [Min21a, Theorem 1.1] and Remark 2.14 that $M$ admits a minimal injective resolution $\left\{\left[M_{i}\xrightarrow{f_{i}}I_{i}\xrightarrow{g_{i}}M_{i-1}\to\right]\right\}_{i\leq 0}$, as in Definition 2.13, such that $I_{-i}=0$ for $i\gg 0$. Let $e\coloneqq\max\{i\text{ : }I_{-i}\neq 0\}$.
Our next goal is—roughly speaking—to show that $R$ admits a $K$-injective resolution $J$ built out of the above minimal injective resolution of $R$. Since the dg-module $R$ satisfies the conditions in [Min21a, Definition 3.1], it follows from [Min21a, Theorem 3.5] that $e=\operatorname{id}_{A}(R)=\dim H^{0}(A)$. We observe that, by [Min21a, Theorem 3.5] and [Min21b, Definition 3.1], we have $\inf(M_{-i})=\inf(I_{-i})=\inf(R)$ for all $i=0,\dots,e$. If $\operatorname{id}_{A}(R)=0$, then $\dim H^{0}(A)=0$, and hence $\dim_{\mathbf{k}}H(A)<\infty$; the statement holds in this case by Proposition 3.15 and Remark 3.17. Assume $\operatorname{id}_{A}(R)\geq 1$. Applying [Min21b, Lemma 3.22] and induction, we have that $\operatorname{id}_{A}(M_{-e})=0$. By [Min21b, Lemma 3.17], we conclude that there is an isomorphism $M_{-e}\cong I[l]$ in $\operatorname{D}(A)$ for some $I\in\mathcal{I}$ and $\ell\in\mathbb{Z}$. By Remark 2.9, $I[l]$ is $K$-injective. Moreover, it follows from [Min21b, Theorem 3.10(4)] that $M_{-e}\to I_{-e}$ is an isomorphism in $\operatorname{D}(A)$, and hence we may assume that $I[l]=I_{e}$. Using the triangle $M_{-e+1}\to I_{-e+1}\to M_{-e}\to$, one may construct a $K$-injective resolution of $M_{-e+1}$ from $I_{-e+1}$ and $I[l]$. Continuing inductively gives a $K$-injective resolution $R\xrightarrow{\simeq}J$ that is built from the data of our minimal injective resolution of $R$.
We have isomorphisms $\mathbf{k}\cong\mathbf{R}\underline{\operatorname{Hom}}_{A}(\mathbf{k},R)\cong\operatorname{\underline{Hom}}_{A}(\mathbf{k},J)$ in $\operatorname{D}(A)$. For all $i\leq 0$, we have isomorphisms $I_{i}\cong\operatorname{\underline{Hom}}_{A^{0}}(A,K_{i})[-\inf(R)]$ in $\operatorname{D}(A)$ for some $K_{i}\in\mathrm{Inj}(A^{0})$. It therefore follows from adjunction that, for all $i\leq 0$, we have $\operatorname{\underline{Hom}}_{A}(\mathbf{k},I_{i})\cong\operatorname{\underline{Hom}}_{A^{0}}(\mathbf{k},K_{i})[-\inf(R)]$ as bigraded complexes of vector spaces. In particular, $\operatorname{\underline{Hom}}_{A}(\mathbf{k},I_{i})$ lives in the single cohomological degree $j\coloneqq\inf(R)$. For all $i,q\in\mathbb{Z}$, we thus have $\operatorname{\underline{Hom}}_{A}(\mathbf{k},I_{i})_{q}\cong\operatorname{Hom}_{\operatorname{D}(A)}(\mathbf{k},I_{i}(q)[j])$. 2.10(1) and [Min21b, Corollary 3.29] imply that $\operatorname{Hom}_{\operatorname{D}(A)}(\mathbf{k},I_{i}(q)[j])\cong\operatorname{Hom}_{H^{0}(A)}(\mathbf{k},H^{0}\operatorname{\underline{Hom}}_{A^{0}}(A,K_{i}(q)))\cong\operatorname{Hom}_{\operatorname{D}(A)}(\mathbf{k},R(q)[j-i])$. This last object is zero unless $q=0$ and $i=j$. If $\operatorname{\underline{Hom}}_{A}(\mathbf{k},I_{i})=0$ for all $i$, it follows that $\operatorname{\underline{Hom}}_{A}(\mathbf{k},J)=0$, a contradiction. Thus, we have $-e\leq j\leq 0$, and $\operatorname{Hom}_{A}(\mathbf{k},J)\cong\operatorname{\underline{Hom}}_{A}(\mathbf{k},I_{j})\cong\mathbf{k}$ in $\operatorname{Mod}(A)$.
It suffices to exhibit an isomorphism $\Gamma_{\mathfrak{m}}(J)\coloneqq\underset{d\to\infty}{\operatorname{colim}}\operatorname{\underline{Hom}}_{A}(A/A_{\geq d},J)\cong\operatorname{\underline{Hom}}_{\mathbf{k}}(A,\mathbf{k})$ as dg-$A$-modules. We note that $\operatorname{\underline{Hom}}_{\mathbf{k}}(A,\mathbf{k})$ is an essential extension of $\mathbf{k}$ (see [Pop73, Section 3.10] for the definition) in $\operatorname{Mod}(A^{\natural})$ and hence also in $\operatorname{Mod}(A)$. By Remark 2.9, $\operatorname{\underline{Hom}}_{\mathbf{k}}(A,\mathbf{k})\cong\operatorname{\underline{Hom}}_{A^{0}}(A,\operatorname{\underline{Hom}}_{\mathbf{k}}(A^{0},\mathbf{k}))\in\mathcal{I}$ is an injective object in $\operatorname{Mod}(A)$. By [Pop73, Lemmas 10.5 and 10.6], $\operatorname{\underline{Hom}}_{\mathbf{k}}(A,\mathbf{k})$ has no proper essential extensions in $\operatorname{Mod}(A)$ and is hence an injective envelope for $\mathbf{k}$ in $\operatorname{Mod}(A)$. The underlying bigraded module of $\Gamma_{\mathfrak{m}}(J)$ coincides with the $\mathfrak{m}$-torsion submodule of $J$ over $A^{\natural}$; thus, $\Gamma_{\mathfrak{m}}(J)$ is essential over $\operatorname{\underline{Hom}}_{A}(\mathbf{k},J)\cong\mathbf{k}$ in $\operatorname{Mod}(A)$. By [Pop73, Proposition 10.7 and Lemma 10.8], there is an injective morphism $f\colon\Gamma_{\mathfrak{m}}(J)\rightarrow\operatorname{\underline{Hom}}_{\mathbf{k}}(A,\mathbf{k})$ in $\operatorname{Mod}(A)$. Since each $I_{i}$ is a bigraded injective $A^{\natural}$-module, so is $\Gamma_{\mathfrak{m}}(I_{j})$. For each $i\neq j$, it follows that $\Gamma_{\mathfrak{m}}(I_{i})=0$, since it is essential over $\operatorname{\underline{Hom}}_{A}(\mathbf{k},I_{i})=0$. We therefore have isomorphisms $\Gamma_{\mathfrak{m}}(J)\cong\Gamma_{\mathfrak{m}}(I_{j})\cong E_{A^{\natural}}(\mathbf{k})\cong\operatorname{\underline{Hom}}_{\mathbf{k}}(A,\mathbf{k})$ in $\operatorname{Mod}(A^{\natural})$. Since each bigraded component of $\operatorname{\underline{Hom}}_{\mathbf{k}}(A,\mathbf{k})$, and hence $\Gamma_{\mathfrak{m}}(J)$, is a finite dimensional $\mathbf{k}$-vector space, it follows that $f\colon\Gamma_{\mathfrak{m}}(J)\to\operatorname{\underline{Hom}}_{\mathbf{k}}(A,\mathbf{k})$ is an isomorphism in $\operatorname{Mod}(A)$. ∎
Remark 3.19.
Let $A$ and $n$ be as in Theorem 3.18. The inequality $j\coloneqq\inf(R)\leq 0$ in the proof of Theorem 3.18 implies that $n\geq\inf(A)$.
Example 3.20.
Theorem 3.18 applies to each of the families of dg-algebras in [BS23, Examples 2.23 - 2.26]. For instance, Theorem 3.18 applies to the generalized Koszul complexes described in [BS23, Example 2.23] and originally studied by Shaul in [Sha21].
Remark 3.21.
Our proof of 3.18 uses in an essential way Minamoto’s structural results for minimal injective resolutions developed in [Min21a]. Since those results require graded commutativity, our argument does as well. A suitable theory of minimal injective resolutions for noncommutative dg-algebras therefore seems to be needed to prove Theorem 3.18 without the commutativity assumption.
We conjecture that the graded commutativity hypothesis can be removed from Theorem 3.18.
Conjecture 3.22.
Let $A$ be a Gorenstein dg-algebra (see Definition 2.7). Choose $a,n\in\mathbb{Z}$ such that there is an isomorphism $\mathbf{R}\underline{\operatorname{Hom}}_{A}(\mathbf{k},A)\cong\mathbf{k}(a)[-n]$ in $\operatorname{D}(A)$. The dg-$A$-$A$-bimodule $R=A(-a)[n]$ is a balanced dualizing dg-module.
A positive answer to Conjecture 3.22 would give a generalization of Yekutieli’s result in [Yek92] that noncommutative Gorenstein $\mathbf{k}$-algebras admit balanced dualizing complexes. As evidence for Conjecture 3.22, apart from Proposition 3.15 and Theorem 3.18 (along with Remark 3.17), we observe that, if $A$ and $R$ are as in Conjecture 3.22, then $\mathbf{R}\Gamma_{\mathfrak{m}}(R)$ and $\underline{\operatorname{Hom}}_{\mathbf{k}}(A,\mathbf{k})$ are quasi-isomorphic complexes of $\mathbf{k}$-vector spaces. Indeed, recall that there is an isomorphism $\mathbf{R}\underline{\operatorname{Hom}}_{A}(\mathbf{k},R)\cong\mathbf{k}$ in $\operatorname{D}(A)$. Applying $\mathbf{R}\underline{\operatorname{Hom}}_{A}(-,A)$ to the short exact sequence $0\to A_{\geq 1}/A_{\geq 2}\to A/A_{\geq 2}\to\mathbf{k}\to 0$ yields that $\underline{\operatorname{Ext}}^{*}_{A}(A/A_{\geq 2},A)$ is $\mathbf{k}$-linearly isomorphic to $H(\underline{\operatorname{Hom}}_{\mathbf{k}}(A,\mathbf{k}))_{0}\oplus H(\underline{\operatorname{Hom}}_{\mathbf{k}}(A,\mathbf{k}))_{-1}$. Continuing in this way, we conclude by induction that $\underline{\operatorname{Ext}}^{*}_{A}(A/A_{\geq d},A)$ is isomorphic as a bigraded $\mathbf{k}$-vector space to $\bigoplus_{i=0}^{d-1}H(\underline{\operatorname{Hom}}_{\mathbf{k}}(A,\mathbf{k}))_{-i}$, and passing to the colimit gives the desired $\mathbf{k}$-linear quasi-isomorphism. The challenge is exhibiting a quasi-isomorphism between $\mathbf{R}\Gamma_{\mathfrak{m}}(R)$ and $\underline{\operatorname{Hom}}_{\mathbf{k}}(A,\mathbf{k})$ as $A$-$A$-bimodules.
4. Serre duality
We now prove our main result, Theorem 1.3, which is a generalization to dg-algebras of Yekutieli-Zhang’s noncommutative Serre Duality Theorem (Theorem 1.2)
4.1. Noncommutative geometry over a dg-algebra
Let $A$ be as in Setup 2.5. As in the introduction (see also [BS23, Section 3]), we let $\operatorname{D}_{\mathrm{Qgr}}(A)$ denote the quotient $\operatorname{D}(A)/\operatorname{D}^{\operatorname{Tors}}(A)$. We also let $\operatorname{D}_{\operatorname{qgr}}(A)$ denote the essential image of the fully faithful embedding $\operatorname{D}^{\operatorname{b}}(A)/\operatorname{D}^{\operatorname{tors}}(A)\hookrightarrow\operatorname{D}_{\mathrm{Qgr}}(A)$. We recall from the introduction that, given $M\in\operatorname{D}(A)$, we let $\widetilde{M}$ denote the corresponding object in $\operatorname{D}_{\mathrm{Qgr}}(A)$. We have canonical triangulated functors
$\Pi:\operatorname{D}(A)\to\operatorname{D}_{\mathrm{Qgr}}(A)\quad\text{and}\quad\pi:\operatorname{D}^{\operatorname{b}}(A)\to\operatorname{D}_{\operatorname{qgr}}(A)$
given by $M\mapsto\widetilde{M}$. If $A=A^{0}$, and $A$ is commutative and standard graded, then $\operatorname{D}_{\mathrm{Qgr}}(A)$ (resp. $\operatorname{D}_{\operatorname{qgr}}(A)$) is equivalent to the derived category of quasi-coherent (resp. coherent) sheaves on $\operatorname{Proj}(A)$.
Remark 4.1.
Suppose $A$ is as in Setup 2.5; in addition, assume $A^{0}$ is commutative and generated in degree 1, and also that $A$ is finitely generated as an $A^{0}$-module. In this case, $A$ determines a sheaf of dg-algebras $\mathcal{A}$ on $\operatorname{Proj}(A^{0})$, and $\operatorname{D}(A)$ (resp. $\operatorname{D}^{\operatorname{b}}(A)$) may be interpreted as the derived category of quasi-coherent (resp. coherent) dg-$\mathcal{A}$-modules: see [BS23, Section 3.2] for details.
The functor $\Pi:\operatorname{D}(A)\to\operatorname{D}_{\mathrm{Qgr}}(A)$ admits a fully faithful right adjoint [BS23, Proposition 3.3], the derived global sections functor, which we denote by $\mathbf{R}\Gamma_{*}$.
Proposition 4.2.
If $A$ is as in Setup 2.5, and $M\in\operatorname{D}(A)$, then there is an isomorphism $\mathbf{R}\Gamma_{*}(\widetilde{M})\cong\underset{d\to\infty}{\operatorname{colim}}\text{ }\mathbf{R}\underline{\operatorname{Hom}}_{A}(A_{\geq d},M)$ in $\operatorname{D}(A)$.
Proof.
Let $M\in\operatorname{D}(A)$. It follows from (the proof of) [Kra10, Proposition 4.9.1] that there is an exact triangle
(4.3) $\mathbf{R}\Gamma_{\mathfrak{m}}(M)\xrightarrow{}M\xrightarrow{}\mathbf{R}\Gamma_{*}(\widetilde{M})\xrightarrow{},$
where the first morphism is the counit of the adjunction between the inclusion $\operatorname{D}^{\operatorname{Tors}}(A)\hookrightarrow\operatorname{D}(A)$ and the functor $\mathbf{R}\Gamma_{\mathfrak{m}}$. On the other hand, for all $d\geq 0$, the exact triangle $A_{\geq d}\to A\to A/A_{\geq d}\to$ induces an exact triangle $\mathbf{R}\underline{\operatorname{Hom}}_{A}(A/A_{\geq d},M)\xrightarrow{}M\xrightarrow{}\mathbf{R}\underline{\operatorname{Hom}}_{A}(A_{\geq d},M)\xrightarrow{};$ passing to colimits, we obtain the exact triangle
(4.4) $\mathbf{R}\Gamma_{\mathfrak{m}}(M)\xrightarrow{}M\xrightarrow{}\underset{d\to\infty}{\operatorname{colim}}\text{ }\mathbf{R}\underline{\operatorname{Hom}}_{A}(A_{\geq d},M)\xrightarrow{}.$
Finally, by Remark 3.5, the first morphisms in the triangles (4.3) and (4.4) coincide. ∎
Recall from the introduction that, given $\widetilde{M},\widetilde{N}\in\operatorname{D}_{\operatorname{Qgr}}(A)$, we write $\operatorname{Ext}^{j}(\widetilde{M},\widetilde{N})\coloneqq\operatorname{Hom}_{\operatorname{D}_{\operatorname{Qgr}}(A)}(\widetilde{M},\widetilde{N}[j])$. Given $A$ as in Setup 2.5 and $M\in\operatorname{Mod}(A)$, let $\mathbf{R}^{j}\Gamma(\widetilde{M})\coloneqq H^{j}\mathbf{R}\Gamma_{*}(\widetilde{M})_{0}$.
Proposition 4.5.
Let $A$ be as in Setup 2.5, $M\in\operatorname{D}^{\operatorname{b}}(A)$, and $N\in\operatorname{D}(A)$. For all $j\in\mathbb{Z}$, we have an isomorphism $\operatorname{Ext}^{j}(\widetilde{M},\widetilde{N})\cong\underset{d\to\infty}{\operatorname{colim}}\operatorname{Ext}^{j}_{A}(M_{\geq d},N).$ In particular, we have $\mathbf{R}^{j}\Gamma(\widetilde{N})\cong\operatorname{Ext}^{j}(\widetilde{A},\widetilde{N})$.
Proof.
The triangle (4.3) and the exactness of filtered colimits gives a long exact sequence
$\cdots\to\underset{d\to\infty}{\operatorname{colim}}\operatorname{Hom}_{\operatorname{D}(A)}(M_{\geq d},\mathbf{R}\Gamma_{\mathfrak{m}}(N)[j])$ $\to\underset{d\to\infty}{\operatorname{colim}}\operatorname{Hom}_{\operatorname{D}(A)}(M_{\geq d},N[j])$
$\to\underset{d\to\infty}{\operatorname{colim}}\operatorname{Hom}_{\operatorname{D}(A)}(M_{\geq d},\mathbf{R}\Gamma_{*}(\widetilde{N})[j])$
$\to\underset{d\to\infty}{\operatorname{colim}}\operatorname{Hom}_{\operatorname{D}(A)}(M_{\geq d},\mathbf{R}\Gamma_{\mathfrak{m}}(N)[j+1])\to\cdots.$
Since $\widetilde{M_{\geq m}}\cong\widetilde{M}$ for all $m\in\mathbb{Z}$, adjunction implies that $\underset{d\to\infty}{\operatorname{colim}}\operatorname{Hom}_{D(A)}(M_{\geq d},\mathbf{R}\Gamma_{*}(\widetilde{N})[j])\cong\operatorname{Ext}^{j}_{A}(\widetilde{M},\widetilde{N}).$ It therefore suffices to show that $\underset{d\to\infty}{\operatorname{colim}}\operatorname{Hom}_{\operatorname{D}(A)}(M_{\geq d},\mathbf{R}\Gamma_{\mathfrak{m}}(N)[j])=0$ for all $j$. Replacing $N$ with an appropriate shift, we may assume without loss that $j=0$. Since $\mathbf{R}\Gamma_{\mathfrak{m}}(N)$ is torsion, it is a filtered colimit of objects in $\operatorname{D}^{\operatorname{tors}}(A)$; thus, since $M\in\operatorname{D}^{\operatorname{b}}(A)$, we need only show $\underset{d\to\infty}{\operatorname{colim}}\operatorname{Hom}_{\operatorname{D}(A)}(M_{\geq d},T)=0$ for $T\in\operatorname{D}^{\operatorname{tors}}(A)$. There is a quasi-isomorphism $T\twoheadrightarrow T/T_{\geq m}$ for $m\gg 0$, and so we may assume $T_{m}=0$ for $m\gg 0$. Let $F^{d}$ be a semifree resolution of $M_{\geq d}$ as in [BS23, Proposition 2.16] for all $d$; the construction of $F^{d}$ shows that $F^{d}_{i}=0$ for $i<d$. We therefore have $\operatorname{Hom}_{\operatorname{D}(A)}(M_{\geq d},T)\cong\operatorname{Hom}_{K(A)}(F^{d},T)=0$ for $d\gg 0$, where the isomorphism follows from Remark 2.2. This proves the first statement, and the second is immediate from the first, along with Proposition 4.2. ∎
4.2. Serre duality
We now prove Theorem 1.3. We begin with the following technical result:
Proposition 4.6.
Let $A$ be a dg-algebra as in Setup 2.5 with a balanced dualizing dg-module $R$, and let $M\in\operatorname{D}^{\operatorname{b}}(A)$. For $j\in\mathbb{Z}$ and $d\geq 1$, we have a canonical isomorphism of $\mathbf{k}$-vector spaces
$\mathbf{R}\Gamma^{j}(\widetilde{M})\cong\operatorname{Hom}_{\mathbf{k}}(\operatorname{Ext}_{A}^{-j-1}(M_{\geq d},R),\mathbf{k}).$
Proof.
Since $\widetilde{M_{\geq d}}=\widetilde{M}$, the triangle (4.3) yields a long exact sequence
$\cdots\to H^{j}(M_{\geq d})_{0}\to\mathbf{R}^{j}\Gamma(\widetilde{M})\to H^{j+1}_{\mathfrak{m}}(M_{\geq d})_{0}\to H^{j+1}(M_{\geq d})_{0}\to\cdots.$
Since $d\geq 1$, the outer terms vanish, yielding the isomorphism $\mathbf{R}^{j}\Gamma(\widetilde{M})\xrightarrow{\cong}H^{j+1}_{\mathfrak{m}}(M_{\geq d})_{0}$. We therefore have isomorphisms
$\mathbf{R}^{j}\Gamma(\widetilde{M})\cong H^{j+1}_{\mathfrak{m}}(M_{\geq d})_{0}\cong\operatorname{Ext}^{j+1}_{A^{\operatorname{op}}}(D(M_{\geq d}),\operatorname{\underline{Hom}}_{\mathbf{k}}(A,\mathbf{k}))$ $\cong\operatorname{Ext}^{j+1}_{\mathbf{k}}(D(M_{\geq d}),\mathbf{k})$
$\cong\operatorname{Hom}_{\mathbf{k}}(\operatorname{Ext}^{-j-1}_{A}(M_{\geq d},R),\mathbf{k}),$
where the second isomorphism follows from Corollary 3.14, the third from adjunction, and the fourth from the exactness of taking $\mathbf{k}$-duals. ∎
Proof of Theorem 1.3.
The canonical isomorphisms $\mathbf{R}^{j}\Gamma(\widetilde{M})\cong\operatorname{Hom}_{\mathbf{k}}(\operatorname{Ext}_{A}^{-j-1}(M_{\geq d},R),\mathbf{k})$ from Proposition 4.6 for $d\geq 1$ are compatible with the maps
$\operatorname{Hom}_{\mathbf{k}}(\operatorname{Ext}_{A}^{-j-1}(M_{\geq d+1},R),\mathbf{k})\to\operatorname{Hom}_{\mathbf{k}}(\operatorname{Ext}_{A}^{-j-1}(M_{\geq d},R),\mathbf{k})$
induced by the inclusions $M_{\geq d+1}\hookrightarrow M_{\geq d}$. Passing to the colimit as $d$ tends to infinity yields the desired isomorphism. ∎
Example 4.7.
If $A$ is a graded commutative and Gorenstein dg-algebra, then Theorem 3.18 implies that $A(-a)[n]$ is a balanced dualizing dg-module for $A$. It therefore follows from Theorem 1.3 that, for all $M\in\operatorname{D}^{\operatorname{b}}(A)$ and $j\in\mathbb{Z}$, we have a $\mathbf{k}$-linear isomorphism
$\mathbf{R}^{j}\Gamma(\widetilde{M})\cong\operatorname{Hom}_{\mathbf{k}}(\operatorname{Ext}^{n-j-1}(\widetilde{M},\widetilde{A(-a)}).$
For instance, this isomorphism holds for each of the families of dg-algebras in [BS23, Examples 2.23—2.26]. A positive answer to Conjecture 3.22 would imply that the graded commutative assumption in this example may be removed.
5. Application to finiteness properties of derived global sections
The goal of this section is to apply our Serre Duality Theorem (Theorem 1.3) to establish some finiteness properties for derived global sections of objects in $\operatorname{D}_{\operatorname{qgr}}(A)$. Let us begin by fixing some notation. Given $i\in\mathbb{Z}$, let $\operatorname{D}(A)_{\geq i}$ denote the subcategory of $\operatorname{D}(A)$ given by objects $M$ such that $M_{j}=0$ for $j<i$, and set $\operatorname{D}^{\operatorname{b}}(A)_{\geq i}\coloneqq\operatorname{D}^{\operatorname{b}}(A)\cap\operatorname{D}(A)_{\geq i}$. The truncation functor $\tau_{\geq i}\colon\operatorname{D}(A)\to\operatorname{D}(A)_{\geq i}$ is the right adjoint to the inclusion map $\operatorname{D}(A)_{\geq i}\hookrightarrow\operatorname{D}(A)$. We conclude that $\mathbf{R}\Gamma_{\geq i}\coloneqq\tau_{\geq i}\circ\mathbf{R}\Gamma_{*}$ is the right adjoint of the composition $\operatorname{D}(A)_{\geq i}\hookrightarrow\operatorname{D}(A)\xrightarrow{\Pi}\operatorname{D}_{\operatorname{Qgr}}(A)$.
Theorem 5.1.
Let $A$ be as in Setup 2.5. Assume $A$ is Gorenstein and that $A$ admits a balanced dualizing dg-module $R$. Let $M\in\operatorname{D}^{\operatorname{b}}(A)$, and fix $i\in\mathbb{Z}$.
  1. (1)
    $H^{j}(\mathbf{R}\Gamma_{\geq i}(\widetilde{M}))$ is a finitely generated $H^{0}(A)$-module for all $j\in\mathbb{Z}$.
  2. (2)
    $\mathbf{R}^{j}\Gamma(\widetilde{M})\neq 0$ only if $\inf(M)\leq j\leq\sup(M)-\inf(R)-1$ (see (2.12) for the definitions of the $\sup$ and $\inf$ of a dg-module).
  3. (3)
[…]
Example 5.2.
Suppose $A$ and $M$ are as in Theorem 5.1. Choose $a,n\in\mathbb{Z}$ such that there is an isomorphism $\mathbf{R}\underline{\operatorname{Hom}}_{A}(\mathbf{k},A)\cong\mathbf{k}(a)[-n]$ in $\operatorname{D}(A)$. If $A$ is graded commutative, then $A(-a)[n]$ is a balanced dualizing dg-module for $A$ by Theorem 3.18. Theorem 5.1(2) therefore implies:
$\mathbf{R}^{j}\Gamma(\widetilde{M})\neq 0\quad\text{only if}\quad\inf(M)\leq j\leq\sup(M)+n-\inf(A)-1.$
Assume, in addition, that $A=A^{0}$, $A$ is generated in internal degree 1, and $M$ is concentrated in degree 0. Let $X=\operatorname{Proj}(A)$. In this case, $n=\dim(A)$, and so Theorem 5.1 recovers the classical statement that the $A$-module $\bigoplus_{\ell\geq i}H^{j}(X,\widetilde{M}(\ell))$ is finitely generated for all $i$, and $H^{j}(X,\widetilde{M})\neq 0$ only if $0\leq j\leq n-1=\dim(X)$.
It was previously known that the functor $\mathbf{R}\Gamma_{\geq i}$ restricts to a functor $\operatorname{D}_{\operatorname{qgr}}(A)\to\operatorname{D}^{\operatorname{b}}(A)_{\geq i}$ for dg-algebras $A$ as in Setup 2.5 in the following cases:
  1. (1)
    A result of Artin-Zhang implies this when $A=A^{0}$, and $A$ satisfies Artin-Zhang’s condition $\chi$ [AZ94, Theorem 3.8(3)].
  2. (2)
    By [BS23, Proposition 3.6], this holds when $A^{0}$ is either Gorenstein or commutative.
As a first step toward proving Theorem 5.1, we formulate a natural extension of Artin-Zhang’s condition $\chi$ to dg-algebras:
Definition 5.3 (cf. [AZ94] Definition 3.7).
Let $A$ be a dg-algebra as in Setup 2.5. We say $A$ satisfies condition $\chi$ if, for all $M\in\operatorname{D}^{\operatorname{b}}(A)$ and $i\in\mathbb{Z}$, we have $\underline{\operatorname{Ext}}^{i}_{A}(\mathbf{k},M)_{j}=0$ for $j\gg 0$.
Suppose $A$ is as in Setup 2.5. If $A=A^{0}$, then Definition 5.3 is equivalent to Artin-Zhang’s condition $\chi$ by [AZ94, Proposition 3.8(1)]. When $A$ is graded commutative, condition $\chi$ always holds: indeed, choosing a semifree resolution $F$ of $\mathbf{k}$ as in [BS23, Proposition 2.16], it is evident that $\operatorname{Ext}^{j}_{A}(\mathbf{k},M)=H^{j}\underline{\operatorname{Hom}}_{A}(F,M)$ is finitely generated over $A^{0}$; since it is also a $\mathbf{k}$-module, condition $\chi$ follows immediately. This argument fails in the noncommutative case, since one cannot always equip the semifree resolution $F$ of $\mathbf{k}$ with a dg-$A$-$A$-bimodule structure. Indeed, there are algebras $A$ as in Setup 2.5 with $A=A^{0}$ such that condition $\chi$ does not hold (see [SZ94, Example 0.1]). The following generalization of a result of Yekutieli-Zhang [YZ97, Corollary 4.3] shows that condition $\chi$ holds for Gorenstein dg-algebras:
Proposition 5.4.
Let $A$ be a dg-algebra as in Setup 2.5. If $A$ is Gorenstein (Definition 2.7), then $A$ satisfies condition $\chi$.
Proof.
Let $M\in\operatorname{D}^{\operatorname{b}}(A)$. Since $A$ is Gorenstein, the natural map
$\mathbf{R}\underline{\operatorname{Hom}}_{A}(\mathbf{k},M)\to\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(\mathbf{R}\underline{\operatorname{Hom}}_{A}(M,A),\mathbf{R}\underline{\operatorname{Hom}}_{A}(\mathbf{k},A))$
is a $\mathbf{k}$-linear quasi-isomorphism. Using again that $A$ is Gorenstein, choose $a,n\in\mathbb{Z}$ such that there is an isomorphism $\mathbf{R}\underline{\operatorname{Hom}}_{A}(\mathbf{k},A)\simeq\mathbf{k}(a)[-n]$ in $\operatorname{D}(A)$. In fact, $\mathbf{R}\underline{\operatorname{Hom}}_{A}(\mathbf{k},A)$ and $\mathbf{k}(a)[-n]$ are isomorphic in $\operatorname{D}(A\otimes_{\mathbf{k}}A^{\operatorname{op}})$, since the quasi-isomorphisms
$\mathbf{R}\underline{\operatorname{Hom}}_{A}(\mathbf{k},A)\xrightarrow{\simeq}\sigma^{\geq n}\mathbf{R}\underline{\operatorname{Hom}}_{A}(\mathbf{k},A)\xleftarrow{\simeq}\sigma^{\leq n}\sigma^{\geq n}\mathbf{R}\underline{\operatorname{Hom}}_{A}(\mathbf{k},A)\cong\mathbf{k}[-n](a)$
are morphisms of dg-$A$-$A$-bimodules; here $\sigma^{\geq n}$ and $\sigma^{\leq n}$ denote smart truncations, as in [BS23, Remark 2.7]. We therefore have a $\mathbf{k}$-linear quasi-isomorphism
$\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(\mathbf{R}\underline{\operatorname{Hom}}_{A}(M,A),\mathbf{R}\underline{\operatorname{Hom}}_{A}(\mathbf{k},A))\simeq\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(\mathbf{R}\underline{\operatorname{Hom}}_{A}(M,A),\mathbf{k}(a)[-n]).$
Since $A$ is Gorenstein, we have $\mathbf{R}\underline{\operatorname{Hom}}_{A}(M,A)\in\operatorname{D}^{\operatorname{b}}(A^{\operatorname{op}})$. Choose a semifree resolution $F$ of the dg-$A^{\operatorname{op}}$-module $\mathbf{R}\underline{\operatorname{Hom}}_{A}(M,A)$ as in [BS23, Proposition 2.16], so that
$\mathbf{R}\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(\mathbf{R}\underline{\operatorname{Hom}}_{A}(M,A),\mathbf{k}(a)[-n])\cong\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(F,\mathbf{k}(a)[-n]).$
Finally, we have $F_{j}=0$ for $j\ll 0$, which implies that $\underline{\operatorname{Hom}}_{A^{\operatorname{op}}}(F,\mathbf{k}(a)[-n])_{j}=0$ for $j\gg 0$. ∎
Before we prove Theorem 5.1, we establish the following two technical results:
Lemma 5.5.
Let $A$ be a dg-algebra as in Setup 2.5, $M\in\operatorname{D}^{\operatorname{b}}(A)$, and $N\in\operatorname{D}(A)$ such that $\inf(N)>-\infty$. We have $\inf(\mathbf{R}\underline{\operatorname{Hom}}_{A}(M,N))\geq\inf(N)-\sup(M)$.
Proof.
Write $\ell\coloneqq\inf(N)$, and let $\sigma^{\geq\ell}N$ denote the smart truncation of $N$ in cohomological degrees at least $\ell$. There is a quasi-isomorphism $N\xrightarrow{\simeq}\sigma^{\geq\ell}N$ of dg-$A$-modules [BS23, Remark 2.7]. Let $F\xrightarrow{\simeq}M$ be a semifree resolution as in [BS23, Proposition 2.16]; it follows from the proof of this result that $F^{i}=0$ for $i>\sup(M)$. We have $\mathbf{R}\underline{\operatorname{Hom}}_{A}(M,N)\cong\underline{\operatorname{Hom}}_{A}(F,\sigma^{\geq\ell}N)$, and the latter complex vanishes in cohomological degrees smaller than $\inf(N)-\sup(M)$. ∎
Proof of Theorem 5.1.
Let $j\in\mathbb{Z}$. Since $H^{j}(M)$ is finitely generated over $H^{0}(A)$, the triangle (4.3) implies that, to prove (1), it suffices to show that $H^{j}_{\mathfrak{m}}(M)_{\geq i}$ is finitely generated over $H^{0}(A)$. We begin by showing that $\dim_{\mathbf{k}}H^{j}_{\mathfrak{m}}(M)_{\ell}<\infty$ for all $\ell\in\mathbb{Z}$. Remark 2.6 implies that $\dim_{\mathbf{k}}H^{t}(M)_{\ell}<\infty$ for all $\ell,t\in\mathbb{Z}$; since $R$ is a dualizing dg-module, $\mathbf{R}\underline{\operatorname{Hom}}(M_{\geq 1},R)\in\operatorname{D}^{\operatorname{b}}(A)$, and so Proposition 4.6 yields that $\dim_{\mathbf{k}}\mathbf{R}^{t}\Gamma(\widetilde{M(\ell)})<\infty$ for all $\ell,t\in\mathbb{Z}$ as well. The triangle (4.3) thus implies that $\dim_{\mathbf{k}}H^{j}_{\mathfrak{m}}(M)_{\ell}<\infty$ for all $\ell\in\mathbb{Z}$. To prove (1), we therefore need only show $H^{j}_{\mathfrak{m}}(M)_{\ell}=0$ for $\ell\gg 0$.
Let $d\in\mathbb{Z}$. The short exact sequence $0\to A_{\geq d}/A_{\geq d+1}\to A/A_{\geq d+1}\to A/A_{\geq d}\to 0$ induces the following exact sequence:
$\operatorname{\underline{Ext}}^{j-1}_{A}(A_{\geq d}/A_{\geq d+1},M)\xrightarrow{}\operatorname{\underline{Ext}}^{j}_{A}(A/A_{\geq d},M)\xrightarrow{}\operatorname{\underline{Ext}}^{j}_{A}(A/A_{\geq d+1},M)\xrightarrow{}\operatorname{\underline{Ext}}^{j}_{A}(A_{\geq d}/A_{\geq d+1},M).$
The dg-$A$-module $A_{\geq d}/A_{\geq d+1}$ is quasi-isomorphic to a finite direct sum of cohomological shifts of $\mathbf{k}(-d)$. Since $A$ satisfies condition $\chi$ (Proposition 5.4), we conclude that there exists $d\gg 0$ such that
$\operatorname{\underline{Ext}}^{j-1}_{A}(A_{\geq m}/A_{\geq m+1},M)_{\geq i}=\operatorname{\underline{Ext}}^{j}_{A}(A_{\geq m}/A_{\geq m+1},M)_{\geq i}=0\quad\text{for all $m\geq d$},$
which implies that the canonical map $\operatorname{\underline{Ext}}^{j}_{A}(A/A_{\geq m},M)_{\geq i}\xrightarrow{}\operatorname{\underline{Ext}}^{j}_{A}(A/A_{\geq m+1},M)_{\geq i}$ is an isomorphism for all $m\geq d$. Thus, $H^{j}_{\mathfrak{m}}(M)_{\geq i}=\operatorname{\underline{Ext}}^{j}_{A}(A/A_{\geq d},M)_{\geq i}$, and so, to prove (1), it suffices to show that $\operatorname{\underline{Ext}}^{j}_{A}(A/A_{\geq d},M)_{\ell}=~{}0$ for $\ell\gg 0$. Since $A/A_{\geq d}\in\operatorname{D}^{\operatorname{tors}}_{\operatorname{gr}}(A)$, and $\operatorname{D}^{\operatorname{tors}}_{\operatorname{gr}}(A)$ is generated by $\mathbf{k}(t)$ for $t\in\mathbb{Z}$ [BS23, Lemma 3.1], this holds since $A$ satisfies condition $\chi$ (Proposition 5.4).
We now prove (2). Recall that $\mathbf{R}^{j}\Gamma(\widetilde{M})=\underset{d\to\infty}{\operatorname{colim}}\operatorname{Ext}^{j}_{A}(A_{\geq d},M)$. By Lemma 5.5 and the exactness of filtered colimits, it follows that $\mathbf{R}^{j}\Gamma(\widetilde{M})=0$ for $j<\inf(M)$. On the other hand, Theorem 1.3 implies that $\mathbf{R}^{j}\Gamma(\widetilde{M})=\underline{\operatorname{Hom}}_{\mathbf{k}}(\underset{d\to\infty}{\operatorname{colim}}\operatorname{Ext}^{-j-1}_{A}(M_{\geq d},R),\mathbf{k})$. Part (2) therefore follows once again from Lemma 5.5 and the exactness of filtered colimits. Part (3) is immediate from (1) and (2). ∎
As an application of Theorem 5.1, we obtain an extension of [BS23, Theorem 1.3], which is a generalization to the setting of dg-algebras of a theorem of Orlov [Orl09, Theorem 2.5]. More precisely, we deduce:
Theorem 5.6.
Let $\mathbf{k}$ be a field and $A$ a dg-algebra as in Setup 2.5. Suppose $A$ is Gorenstein (Definition 2.7) with Gorenstein parameter $a$, and assume $A$ admits a balanced dualizing dg-module (3.13). Write $\operatorname{D}_{\operatorname{sg}}(A)\coloneqq\operatorname{D}^{\operatorname{b}}(A)/\operatorname{Perf}(A)$, where $\operatorname{Perf}(A)$ denotes the thick subcategory of $\operatorname{D}^{\operatorname{b}}(A)$ generated by $A(i)$ for all $i\in\mathbb{Z}$. Let $q:\operatorname{D}^{\operatorname{b}}(A)\to\operatorname{D}_{\operatorname{sg}}(A)$ and $\pi:\operatorname{D}^{\operatorname{b}}(A)\to\operatorname{D}_{\operatorname{qgr}}(A)$ denote the canonical functors. The objects $\pi A(j)\in\operatorname{D}_{\operatorname{qgr}}(A)$ and $q\mathbf{k}(j)\in\operatorname{D}_{\operatorname{sg}}(A)$ are exceptional for all $j\in\mathbb{Z}$, and we have:
  1. (1)
    If $a>0$, then for each $i\in\mathbb{Z}$, there is a fully faithful functor $\Phi_{i}:\operatorname{D}_{\operatorname{sg}}(A)\to\operatorname{D}_{\operatorname{qgr}}(A)$ and a semiorthogonal decomposition $\operatorname{D}_{\operatorname{qgr}}(A)=\langle\pi A(-i-a+1),\dots,\pi A(-i),\Phi_{i}\operatorname{D}_{\operatorname{sg}}(A)\rangle.$
  2. (2)
    If $a<0$, then for each $i\in\mathbb{Z}$, there is a fully faithful functor $\Psi_{i}:\operatorname{D}_{\operatorname{qgr}}(A)\to\operatorname{D}_{\operatorname{sg}}(A)$ and a semiorthogonal decomposition $\operatorname{D}_{\operatorname{sg}}(A)=\langle q\mathbf{k}(-i),\dots,q\mathbf{k}(-i+a+1),\Psi_{i}\operatorname{D}_{\operatorname{qgr}}(A)\rangle.$
  3. (3)
    If $a=0$, then there is an equivalence $\operatorname{D}_{\operatorname{sg}}(A)\xrightarrow{\simeq}\operatorname{D}_{\operatorname{qgr}}(A)$.
Proof.
Theorem 5.1(3) provides an analogue of [BS23, Proposition 3.6] in our setting; with this in hand, the proof of 5.6 is exactly the same as that of [BS23, Theorem 1.3]. ∎
Remark 5.7.
A positive answer to Conjecture 3.22 would allow one to remove from Theorems 5.1 and 5.6 the assumption of the existence of a balanced dualizing dg-module.